OpenStudy (anonymous):

Simplify: (3x2 − 2) + (5x2 + 5x − 1) 8x2 + 5x − 3 8x2 + 5x − 1 2x2 + 5x − 3 8x2 + 7x − 1

OpenStudy (anonymous):

Combine like terms.

OpenStudy (anonymous):

10x?

OpenStudy (anonymous):

How did you get 10x?

OpenStudy (anonymous):

When you combine like terms, you just combine the numbers that has the same variable and/or exponents after them,

OpenStudy (anonymous):

I did 3x+5x+5x=13x right?

OpenStudy (anonymous):

But weren't there exponents after the 3x and 5x? ("3x2" − 2) + ("5x2" + 5x − 1)

HanAkoSolo (jamierox4ev3r):

it helps to pretend that both of the equations are stacked upon each other and you are basically just adding (like terms)

OpenStudy (anonymous):

I'm Confused :(

HanAkoSolo (jamierox4ev3r):

(3x2 − 2) + (5x2 + 5x − 1) +5x2+5x-1 8x2+5x-3

HanAkoSolo (jamierox4ev3r):

it is merely basic arithmetic

HanAkoSolo (jamierox4ev3r):

3x^2+5x^2=8x2 5x=5x -2+(-1)=-3

HanAkoSolo (jamierox4ev3r):

when you add all of these together, your simplified product is 8x^2=5x-3 (standard form)

HanAkoSolo (jamierox4ev3r):

does that make sense @SOS101 ?

OpenStudy (anonymous):

That makes a lot more sense, Thank you very much!

HanAkoSolo (jamierox4ev3r):

no problem XD